OpenStudy (anonymous):

Convert the coordinates to polar cordinates: (1,1)

OpenStudy (anonymous):

nkili after you help her can you help me on my latest q?

OpenStudy (anonymous):

( \[(\sqrt{2}, \pi/4)\]

OpenStudy (anonymous):

Would you like it explained?

OpenStudy (anonymous):

duke can u help with my q?

OpenStudy (anonymous):

1^2 + 1^2 = r^2 r^2 = 2 \[r=\sqrt{2}\] \[\theta=\tan^{-1} (1)=pi/4\] Since a polar coordinate is (r, theta), it is \[(\sqrt{2}, \pi/4)\]

OpenStudy (anonymous):

what question?